[Brahms-l] Shift report 20050529 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun May 29 2005 - 00:00:24 EDT
Shifter leader was rk
Shifters were ps


1. Quite good, some problems at the beginning of the shift with keeping the beam. Since then OK.
2. Data runs: 15026 15027 15030 15031 15032 15033 15034 15035
3. BB3 mainframe is not working as usual, discovered T4 problems with 2U planes, every even wire above 14 has low counts. Reseting low voltage did not work - expert will look at this tomorrow.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun May 29 00:00:40 2005

This archive was generated by hypermail 2.1.8 : Sun May 29 2005 - 00:00:52 EDT